# The Higgs Mass and a Perspective on Broken Autonomy of Scales

### v1

@@
Abstract.
The hierarchy problem of the Higgs mass, the violation of the autonomy of far-distant energy scales, is identified as a pseudo-problem. The pseudo-problem is based on the dogmatic adherence to the methodology of effective theories, for which there is no justification when dealing with presumably fundamental phenomena such as the Higgs mechanism. In view of further breaches of autonomy of scales in fundamental principles outside particle physics, the hierarchy problem is instead reinterpreted as an indication of the fundamental ontological status of the Higgs boson. Selective Realism justifies this attribution within the effective theories of the standard model of particle physics. Moreover, a model of changes in aesthetic canons provides an explanation for the existing dogmatism regarding effective theories.
Key words and phrases:
@@ 1. Introduction
This article claims that the restoration of technical naturalness of the Higgs mass with new physics represents an engagement with a pseudo-problem, thus hindering the progress of the BSM physics program. We will present arguments that there is no need for a physical explanation of a violation of technical naturalness in this particular case. This far-reaching critique of the presumably most crucial non-empirical guiding principle of BSM physics, which significantly contributed to the multibillion-dollar construction of the LHC in Geneva [^2] Of course, confirming the hypothetical Higgs boson itself was perhaps an even stronger reason to build the LHC. However, this discovery was not its only purpose., is not isolated. In recent times, several calls for a re-evaluation of naturalness have been made. Prominent critics include the CERN theorist Giudice (2017) and the author, physicist, and philosopher Hossenfelder (among others: 2018, 2019). While Giudice, formerly a staunch defender of naturalness, offers mild criticism and proclaims a post-naturalness era but wishes to retain naturalness as a still significant criterion, Hossenfelder is uncompromising. She calls for a complete departure from the principle of naturalness, which she believes has misled particle physics for too long [^3] Other contributions to the naturalness debate include (Giudice 2008, Wells 2013 Dine 2015, Hossenfelder 2017, Williams 2015, 2019, Wallace 2019, Bain 2019, Franklin 2020, Fischer 2020, Koren 2020, Athanasiou 2022). Our apologies go to all authors whose contributions were not mentioned..
@@
In this article, at least some significance should be attributed to the value of the Higgs mass, not creating a need for explanation, but suggesting that the Higgs boson is ontologically fundamental and placing it outside the provisional, effective concepts of the standard model physics. The breakdown of autonomy of scales marks the end of an instrumentalist, pragmatic approach in particle physics that cannot meet the claim of fundamental physics to make definitive statements about the world - ultimately, the desire to relieve philosophy of the task of ontology. In short, instead of being horrified by the failure of effective theories, a sensitivity to large energy scales should provide ample reason for optimism that the overarching goal of fundamental physics will succeed.
@@
The argumentation will proceed as follows: In Chapter 2, we first recapitulate some important definitions of naturalness and particularly describe the lack of technical naturalness of the renormalized Higgs mass. Confronted with this problem, there are three options: firstly, attempting to address it with new physics but repeatedly facing disappointment due to the experimental failures of all previous theories. Secondly, attempting to completely deny the problem: the Higgs mass is a constant of nature, and discussing its value becomes unnecessary. We opt for the third option and question the adherence to effective field theories, but derive a positive reaction to the violated autonomy of scales. To do so, we argue in Chapter 3 that quantum corrections to the Higgs mass can indeed be ascribed a physical reality, and the dependence of the Higgs mass on high-energy domains has significance for theory evaluation: the breakdown of autonomy of scales. However, we then narrow down the scope of autonomy of scales by explaining several phenomena in which its violation occurs. For each of these phenomena, a fundamental ontological status can be attributed, and the violation of autonomy of scales does not pose a problem. In Chapter 4, we argue against the dogmatic adherence to this principle and thus to effective theories as the only way to do physics, as the philosophical consequences of this adherence are incoherent with the actual modus operandi of physics and the worldview revealed therein. Finally, with Selective Realism, we attempt to reconcile the fundamental nature of the Higgs mass with the effective theories of the standard model. Chapter 5 aims to provide the philosophical framework for (irrational) dogmatic adherence to autonomy of scales through McAllister’s theory of scientific change by the change of aesthetic ideals. In the conclusion of Chapter 6, the significant role of effective theories for a thriving future particle physics is emphasized. Nonetheless, occasional detachment from their principles is necessary to avoid pseudo-problems like that of the unnatural Higgs mass.
2. Technical Naturalness and the Hierarchy Problem
2.1. Definitions of Naturalness
@@
In the context of the standard model, to the best of my knowledge, Weinberg (1972) first mentions naturalness in an attempt to find a "natural explanation of the approximate symmetries in nature". Symmetry, whose profound significance for particle physics is undoubted, can restrict the demand for absolute naturalness. This ’t Hooft naturalness (1979) allows $c_{0}$ to take a much smaller value than $\mathcal{O}(1)$ as long as the limit $c_{0}\to 0$ leads to a symmetry. Chiral symmetries or gauge symmetries, for example, protect the fermion masses or gauge boson masses of the standard model from (unnatural) large correction terms in interaction with virtual particles, justifying the constraint imposed by ’t Hooft. Mathematically, this cannot happen for the Higgs boson since this particle, being the only one, has no spin - its mass is ’t Hooft-unnatural. The definition of naturalness that is more relevant for this chapter is reproduced from the ’t Hooftian perspective but further narrows down the set of unnatural phenomena. In this work, this is defined as technical naturalness [^4] In the literature, this term is sometimes used synonymously with ’t Hooft naturalness. Instead, in the present context, the term AoS naturalness (autonomy of scales) is used. However, the unfortunate term technical naturalness, sounding somewhat contradictory, cannot be avoided even in this convention.. It describes the prohibition of mutual influence from energy scales separated by many orders of magnitude (autonomy of scales) and first appears in Susskind (1979). We will sharpen that definition after having provided somewhat more background.
@@
2.2. Renormalization of the Higgs Mass
@@
A clear distinction decoupling and technical naturalness has already been attempted. Franklin (2020) differentiates between autonomy of microstates (IR dynamics is autonomous despite changes in the UV range) and autonomy of microlaws (IR dynamics is autonomous despite changes in UV parameters) [^5] In a very similar way, Bain (2019) calls this difference heuristic and precise decoupling.. However, for the lessons learned from the subsequent argumentation, it will be more helpful to differentiate between weak and strong autonomy of scales: Weak autonomy guarantees that an effective theory $T^{eff}_{i}$, defined on an interval $[L_{i},\Lambda_{i}]$ (IR cutoff $L_{i}$ and UV cutoff $\Lambda_{i}$) empirically adequately describes its intended phenomena, i.e., on intervals $[0,L_{i}]$ and especially $[\Lambda_{i},\infty)$, there are no crucial additional effects that could seriously influence the behavior of the phenomena. In other words, it concerns the real correspondence to distant scales. This weak autonomy legitimizes the application of EFT. In the case of strong autonomy, contributions to phenomena in $T^{eff}_{i}$, expressed in units of $\Lambda_{i}$, are roughly of the order of $\mathcal{O}(10^{1})$. However, their violation does not prevent the application of EFT due to the option of renormalization. The quantum corrections are only a crucial additional effect in the non-renormalized equations, which is then mathematically manageable. In other words, it concerns the virtual correspondence to distant scales.
An analogy may help: A researcher is confronted with a comprehensive textbook on the knowledge of an entire discipline. They want to learn everything important about a specific topic in this discipline and read the book effectively: In the table of contents they find that the topic is covered in Chapter 3, so they focus on this section, and ignore the rest of the book. Weak autonomy is violated if crucial information on the subject can be found in Chapter 28, which cannot be neglected to understand it. The researcher would now have to make significant efforts and read the entire book. Strong autonomy is already violated if, for example, a footnote in Chapter 3 refers to additional content in Chapter 28. Although the content is artificially brought into the limited section of Chapter 3 via the footnote, and effective reading is not hindered, suspicion arises: The topic could have greater importance than initially assumed, could possibly occur universally and fundamentally, if there are correspondences to completely different sections of the book.
@@ 2.3. How to Not Solve the Problem
  • The calculation of corrections in dimensional regularization makes the cutoff disappear. Yes, but the corrections will also take a huge value, as now the heaviest fermion mass $M^{2}$ of the effective theory takes its role (Williams 2015).
@@
  • Supersymmetry provides scalar loop corrections that could restore naturalness. Yes, but the crucial regions of the parameter space are already excluded. Unexcluded values for loop corrections may lower the value of the corrections, but the theory remains unnatural to some degree (Giudice 2017).
@@
Finally, let’s consider the discussion around probability distributions of parameters diagnosed with unnaturalness. The absence of naturalness is often perceived as a problem because small quantities that do not allow variation seem improbable. As emphasized regularly in the literature (Hossenfelder 2019, Williams 2019), however, such a statement requires a probability distribution, which typically implies a multiverse, where, according to that distribution, all possible parameter values are realized in different universes. Otherwise, the attribution of "improbable" would be an unphysical statement, nothing more than a disturbance of aesthetic perception without consequences for the validity of the theory. Due to the ignorance about the shape of such a distribution, if it exists, the demand for naturalness should be abandoned.
Including a probability distribution and the multiverse in the discussion about naturalness seems, however, unhelpful. Firstly, from a pragmatic viewpoint, the naturalness debate does not gain higher physical content with the addition of a multiverse. The concept of a multiverse in the sense of string theory [^6] The multiverse needed for the argumentation here is clearly distinguished from the multiverse of Everett’s many-worlds interpretation of quantum mechanics (Everett 1957). In Tegmark’s (2008) definition, we are dealing with a Level II multiverse in this case. remains a purely mathematical concept that evades any empirical verification and presumably always will. But let’s assume that one could derive a probability distribution of the relevant parameter values and empirically test it. Shifting the discussion about specific values in nature into the multiverse would then represent a complicating relocation of the problem without obtaining insight, as we shall argue in the following. No matter how probable the discovery of values in our universe is, the likelihood is not zero by construction of the distribution. After all, the values are realized in at least one universe, being ours. Turning the naturalness debate into a discussion of a probability distribution of parameters in the multiverse is intended to answer the question of whether we live in a probable realization of SM parameters - hence a natural world without the need for explanation. However, due to the concrete realization of any parameter value according to a given distribution in some universe, the need for explanation is equally absent in every conceivable set of parameters. Noticing that our universe represents an improbable parameter constellation would not be a problem but an unassessable fact. Knowledge of a distribution is therefore irrelevant. With or without the multiverse, the tautological statement remains: The parameters are what they are. It is inconsequential whether they are unchangeable (universe) or variable and possibly improbable (multiverse) - in both options, they cannot be called unnatural.
3. Autonomy of Scales in Effective and Fundamental Theories
Confronted with the experimental value of the Higgs mass, one initially has several options to respond to the dependence of perturbative corrections on physics at extremely high energies: The BSM enthusiast suspects new physics and is disappointed by the experiments of the past decades (Giudice 2017). I’m not saying these experiments were in vain. If a single SM particle evades the otherwise reliable decoupling of low-energy scales from UV physics, this may have an explanation within particle physics. Temporarily ruling this out can and should be considered asa non-negotiable part of scientific practice. The disappointed one can now simplify things and claim: Only the measured physical value of the Higgs mass has significance for the theory, a constant of nature: It is what it is, and any discussion about its numerical value is unnecessary (like Hossenfelder 2018). This would be a prime example of a pseudo-problem. However, as we will see below, this other extreme cannot be an appropriate answer. So, the third option remains to break out of internal physical reasoning and question the methodology of theoretical particle physics (the exclusive use of effective theories) itself. This has already happened in the recent past, albeit with sometimes drastic consequences for physics. Wallace (2019), for instance, warns that the rejection of the principle of naturalness implies a rejection of reductionism as a overarching principle of foundational research. In the following, however, a positive consequence of rejecting the principle of naturalness will be argued for.
3.1. Bare Masses and Loop Corrections
@@
Firstly, it is decisive to clarify the physical reality of the various elements of the renormalization process in order to argue against the second option. In characterizing the hierarchy problem, Wetterich (1984, p. 217) writes in one of the earliest analyses: "In any case, fine-tuning of bare parameters is not really the relevant problem (…) The relevant parameters are the physical parameters, since any predictions of a model must finally be expressed in terms of these". It is, therefore, unproblematic to compensate for the enormous correction terms $\propto\Lambda^{2}$ by an appropriate bare mass; in the end, the renormalized mass remains as the only isolated measurable quantity. Whether the correction terms have a dependence $\propto\log(\Lambda)$ or $\propto\Lambda^{2}$, there is no problematic fine-tuning, and the hierarchy problem loses significance. This part of Wetterich’s analysis of the hierarchy problem is undoubtedly correct.
@@
3.2. The Applicability of Technical Naturalness
With these foundations, we can turn to a critical examination of the scope of autonomy of scales in fundamental physics. Consider the impact of the UV-IR dialogue occurring in the context of the Higgs mass on the widely adopted methodology of EFTs. In the tower of quasi-autonomous domains in the sense of Cao and Schweber (1993), each described by an EFT, the existence of a scalar elementary particle breaks this autonomy and establishes a dialogue between significantly different energy scales. Is it a coincidence that this phenomenon appears in several mechanisms in the universe, which can be considered fundamental for good reasons?
In three mechanisms, such a delicate dependence of large length scales or small energy scales on the other end of the spectrum occurs. Firstly, the specific properties of the quantum vacuum at the smallest length scales influence the expansion rate of the entire universe (vacuum energy), which undoubtedly plays out on the largest available length scales when calculating the cosmological constant (see also Craig 2022 for an overview). Unlike the balancing act performed between the electroweak scale ($10^{-18}$ m) and the Planck scale ($10^{-35}$ m) in the hierarchy problem of the Higgs mass, quantum effects in cosmology are responsible for considerably larger length scales. On the other hand, effective theories usually separate cosmology and QFT999 This passage only addressed the principal influence of the quantum vacuum on the largest possible length scales. Moreover, the violated strong autonomy of scales exists in a way mathematically analogous to the Higgs mass: Summing the bare cosmological constant from general relativity and the value of the vacuum energy, $\tilde{\lambda}+\lambda=\Lambda$, results in handling a mismatch of 120 orders of magnitude between quantum effects and the observable value $\Lambda$ (to be distinguished from the cutoff)..

### v2

# Deriving Ontological Statements from the Unnatural Higgs Mass

@@
Abstract
We provide novel, metatheoretical arguments strengthening the position that the naturalness problem of the light Higgs mass is a pseudo-problem: No physics beyond the standard model of particle physics is needed to explain the small value of the Higgs boson. By evaluating previous successes of the guiding principle of technical naturalness, we restrict its applicability to non-fundamental phenomena in the realm of provisional theories within limited energy scales. In view of further breaches of autonomy of scales in apparently fundamental phenomena outside particle physics, the hierarchy problem of the Higgs mass is instead reinterpreted as an indication of the ontologically fundamental status of the Higgs boson. Applying the concept of Selective Realism justifies this seemingly contradictory attribution within the effective theories of the standard model of particle physics. Moreover, we argue that the ongoing naturalness debate about the Higgs mass is partly based on the adherence to the methodology of effective theories (often claimed to be universally applicable), for which there is no justification when dealing with presumably fundamental phenomena such as the Higgs mechanism, even if it is embedded into an effective theory.
keywords:
@@ 1 Introduction
This article provides novel (metatheoretical) arguments that the restoration of technical naturalness of the Higgs mass with new physics represents an engagement with a pseudo-problem, thus hindering the progress of the BSM physics program. We will discuss reasons that there is no need for a physical explanation of a violation of technical naturalness in this particular case. This critique of the presumably most crucial non-empirical guiding principle of BSM physics, which significantly contributed to the multibillion-dollar construction of the LHC in Geneva [^2] Of course, confirming the hypothetical Higgs boson itself was perhaps an even stronger reason to build the LHC. However, this discovery was not its only purpose., is not isolated. In recent times, several calls for a re-evaluation of naturalness have been made. Prominent critics include the CERN theorist Giudice (2017) and the author, physicist, and philosopher Hossenfelder (among others: 2018, 2021). While Giudice, formerly a staunch defender of naturalness, offers mild criticism and proclaims a post-naturalness era but wishes to retain naturalness as a still significant criterion, Hossenfelder is uncompromising. She calls for a complete departure from the principle of naturalness, which she believes has misled particle physics for too long [^3] Other contributions to the naturalness debate include (Giudice 2008, Wells 2013 Dine 2015, Hossenfelder 2017, Williams 2015, 2019, Borrelli and Castellani 2019, Wallace 2019, Bain 2019, Harlander and Rosaler 2019, Franklin 2020, Koren 2020, Rosaler 2021, Athanasiou 2022, Fischer 2023, 2024). Our apologies go to all authors whose contributions were not mentioned..
@@
In this article, at least some significance - in contrast to Hossenfelder - should be attributed to the small value of the Higgs mass, not creating a need for explanation, but indicating that the Higgs boson is ontologically fundamental and placing it outside the provisional, effective concepts of the standard model physics. The unnatural Higgs mass does not pose a problem for high-energy physics, but instead allows for ontological commitments about high-energy physics. The breakdown of autonomy of scales marks the end of an instrumentalist, pragmatic approach in particle physics that cannot meet the claim of fundamental physics to make definitive statements about the world - ultimately, the desire to relieve philosophy of the task of ontology. Instead of being horrified by the failure of effective theories, a sensitivity to large energy scales should provide ample reason for optimism that the overarching goal of fundamental physics will succeed. In summary, the following consecutive hypotheses will be developed:
$H_{1}$: The metatheoretical guiding principle of technical naturalness is not universally applicable, but can only be successfully applied to non-fundamental phenomena modelled in provisional (because instrumentalist), effective theories.
Applying that to the Higgs boson renews the debate about a hierarchy pseudo-problem, adding some novel philosophical arguments: Making use of $H_{1}$ and claiming ontological fundamentality of the Higgs boson, it would not, by definition, have a technical naturalness problem. Thus, we will go one step further and conclude a second hypothesis:
$H_{2}$: Violated autonomy of scales in case of the Higgs is rather an indication of its fundamental ontological status and its special role outside of effective theories. The Higgs boson is thus a realistic element embedded in a mostly instrumentalist theory, the Standard Model.
@@
The argumentation for these hypotheses will proceed as follows: In Chapter 2, we first recapitulate some important definitions of naturalness and particularly describe the lack of technical naturalness of the renormalized Higgs mass. Confronted with this problem, there are three options: firstly, attempting to address it with new physics but repeatedly facing disappointment due to the experimental failures of all previous theories. Secondly, attempting to completely deny the problem: the Higgs mass is a constant of nature, and discussing its value becomes unnecessary. We opt for the third option and question the adherence to effective field theories. To do so, we argue in Chapter 3 that quantum corrections to the Higgs mass can indeed be ascribed a physical reality, and the dependence of the Higgs mass on high-energy domains has significance for theory evaluation: the breakdown of autonomy of scales. However, we then narrow down the scope of autonomy of scales by explaining several phenomena in which its violation occurs. For each of these phenomena, a fundamental ontological status can be attributed, and the violation of autonomy of scales does not pose a problem ($H_{1}$). In Chapter 4, we argue against the dogmatic adherence to this principle and thus to effective theories as the only way to do physics, as the philosophical consequences of this adherence are incoherent with the actual modus operandi of physics and the worldview revealed therein. Finally, with Selective Realism, we attempt to reconcile the fundamental nature of the Higgs mass with the effective theories of the standard model diagnosing a pseudo-problem of the small Higgs mass by means of $H_{2}$. Hence, we derive a positive reaction to the violated autonomy of scales in case of the Higgs boson - that humankind uncovered a fundamental entity and did not created a preliminary instrumentalist construction. In the conclusion of Chapter 5, the significant role of effective theories for a thriving future particle physics is emphasized. Nonetheless, occasional detachment from their principles is necessary to avoid pseudo-problems like that of the unnatural Higgs mass.
2 Technical Naturalness and the Hierarchy Problem
2.1 Definitions of Naturalness
@@
In the context of the standard model, to the best of our knowledge, Weinberg (1972) first mentions naturalness in an attempt to find a "natural explanation of the approximate symmetries in nature". Symmetry, whose profound significance for particle physics is undoubted, can restrict the demand for absolute naturalness. This ’t Hooft naturalness (1979) allows $c_{0}$ to take a much smaller value than $\mathcal{O}(1)$ as long as the limit $c_{0}\to 0$ leads to a symmetry. Chiral symmetries or gauge symmetries, for example, protect the fermion masses or gauge boson masses of the standard model from (unnatural) large correction terms in interaction with virtual particles, justifying the constraint imposed by ’t Hooft. Mathematically, this cannot happen for the Higgs boson since this particle, being the only one, has no spin - its mass is ’t Hooft-unnatural. The definition of naturalness that is more relevant for this chapter is reproduced from the ’t Hooftian perspective but further narrows down the set of unnatural phenomena. In this work, this is defined as technical naturalness [^4] In the literature, this term is sometimes used synonymously with ’t Hooft naturalness. Instead, in the present context, the term AoS naturalness (autonomy of scales) is used. However, the unfortunate term technical naturalness, sounding somewhat contradictory, cannot be avoided even in this convention.. It describes the prohibition of mutual influence from energy scales separated by many orders of magnitude (autonomy of scales) and first appears in Susskind (1979). We will sharpen that definition after having provided somewhat more background [^5] Readers interested in further details about the history of the naturalness debate will find a very readable introduction in (Borrelli, Castellani 2019)..
@@
2.2 Renormalization of the Higgs Mass
@@
A clear distinction between decoupling and technical naturalness has already been attempted. Franklin (2020) differentiates between autonomy of microstates (IR dynamics is autonomous despite changes in the UV range) and autonomy of microlaws (IR dynamics is autonomous despite changes in UV parameters) [^6] In a very similar way, Bain (2019) calls this difference heuristic and precise decoupling.. However, for the lessons learned from the subsequent argumentation, it will be more helpful to differentiate between weak and strong autonomy of scales:
Weak autonomy of scales guarantees that an effective theory $T^{eff}_{i}$, defined on an interval $[L_{i},\Lambda_{i}]$ (IR cutoff $L_{i}$ and UV cutoff $\Lambda_{i}$) empirically adequately describes its intended phenomena, i.e., on intervals $[0,L_{i}]$ and especially $[\Lambda_{i},\infty)$, there are no crucial additional effects that could seriously influence the behavior of the phenomena. In other words, it concerns the real correspondence to distant scales. This weak autonomy legitimizes the application of EFT.
In the case of strong autonomy of scales, contributions to phenomena in $T^{eff}_{i}$, expressed in units of $\Lambda_{i}$, are roughly of the order of $\mathcal{O}(10^{1})$. However, their violation does not prevent the application of EFT due to the option of renormalization. The quantum corrections are only a crucial additional effect in the non-renormalized equations, which is then mathematically manageable. In other words, it concerns the virtual correspondence to distant scales.
An analogy may help to digest the technical definition: A researcher is confronted with a comprehensive textbook on the knowledge of an entire discipline. They want to learn everything important about a specific topic in this discipline and read the book effectively: In the table of contents they find that the topic is covered in Chapter 3, so they focus on this section, and ignore the rest of the book. Weak autonomy is violated if crucial information on the subject can be found in Chapter 28, which cannot be neglected to understand it. The researcher would now have to make significant efforts and read the entire book. Strong autonomy is already violated if, for example, a footnote in Chapter 3 refers to additional content in Chapter 28. Although the content is artificially brought into the limited section of Chapter 3 via the footnote, and effective reading is not hindered, suspicion arises: The topic could have greater importance than initially assumed, could possibly occur universally and fundamentally, if there are correspondences to completely different sections of the book.
@@ 2.3 How to Not Solve the Problem
  • The calculation of corrections in dimensional regularization makes the cutoff disappear. Yes, but the naturalness problem will enter again through the backdoor, as worked out, e.g., by Williams (2015) [^7] For the interested reader experienced in theoretical physics, there are more technical publications like (Harlander and Rosaler 2019a, 2019b) and (Rosaler 2021) discussing several naturalness formulations using a regulator-independent parametrization as well as various concepts of cutoffs and renormalization. We restrict ourselves to a non-technical, but philosophical (metatheoretical) perspective in this article, hopefully providing new insights at a different frontier.
  • Supersymmetry provides scalar loop corrections that could restore naturalness. Yes, but the crucial regions of the parameter space are already excluded. Unexcluded values for loop corrections may lower the value of the corrections, but the theory remains unnatural to some degree (Giudice 2017, Fischer 2024).
@@
3 Autonomy of Scales in Effective and Fundamental Theories
Confronted with the experimental value of the Higgs mass, one initially has several options to respond to the dependence of perturbative corrections on physics at extremely high energies that go beyond the explanations ruled out in Section 2.3:
  1. a)
    Physics beyond the standard model will some day provide a physical explanation for the light Higgs mass in some future (physical explanation).
  2. b)
    Only the measured physical value of the Higgs mass has significance for the theory, being a constant of nature, and any discussion about its numerical value is unfruitful (no explanation needed at all).
  3. c)
    One needs to break out of internal physical reasoning and question the methodology of theoretical particle physics (especially the universal trust in effective field theories) itself (metatheoretical explanation).
The BSM enthusiast in a) suspects new physics and is disappointed by the experiments of the past decades (Giudice 2017). I’m not saying these experiments were in vain. If a single SM particle evades the otherwise reliable decoupling of low-energy scales from UV physics, this may have an explanation within particle physics. Temporarily ruling this out can and should be considered as a non-negotiable part of scientific practice. Option a) is becoming less likely from year to year. After decades of missing experimental evidence, we claim its unplausibility at least within the scope of this article. Instead, we turn to options b) and c), which have philosophical points of contact.
The disappointed BSM enthusiast can simplify things and propose explanation b). The Higgs mass is what it is, and its numerical value should not be questioned at all (like Hossenfelder 2018): a prime example of a pseudo-problem. However, as we will explain, this other extreme cannot be an appropriate answer - a new ontological perspective might help to exclude b). Afterwards, we mainly focus on the metatheoretical treatment of naturalness problems as indicated in the option c). This has already happened in the recent past, albeit with sometimes drastic, consequences for physics and metaphysics. Wallace (2019), for instance, warns that the rejection of the principle of naturalness implies a rejection of reductionism as a overarching principle of foundational research. In this article, however, a positive consequence of rejecting the principle of naturalness will be argued for, shedding an encouraging light on the naturalness debate about the Higgs mass (condensed in $H_{2}$).
3.1 An Ontology of Bare Masses and Loop Corrections
@@
Before turning to response c) , it is decisive to clarify the physical reality of the various elements of the renormalization process in order to argue against option b). In characterizing the hierarchy problem, Wetterich (1984, p. 217) writes in one of the earliest analyses: "In any case, fine-tuning of bare parameters is not really the relevant problem (…) The relevant parameters are the physical parameters, since any predictions of a model must finally be expressed in terms of these". It is, therefore, unproblematic to compensate for the enormous correction terms $\propto\Lambda^{2}$ by an appropriate bare mass; in the end, the renormalized mass remains as the only isolated measurable quantity. Whether the correction terms have a dependence $\propto\log(\Lambda)$ or $\propto\Lambda^{2}$, there is no problematic fine-tuning, and the hierarchy problem loses significance. This part of Wetterich’s analysis of the hierarchy problem is undoubtedly correct.
@@
3.2 Limited Applicability of Technical Naturalness?
With these foundations, we can turn to a critical examination of the scope of autonomy of scales in fundamental physics, thereby elaborating reaction c) to the hierarchy problem. Consider the impact of the UV-IR dialogue occurring in the context of the Higgs mass on the widely adopted methodology of EFTs. In the tower of quasi-autonomous domains in the sense of Cao and Schweber (1993), each described by an EFT, the existence of a scalar elementary particle (the Higgs) breaks this autonomy and establishes a dialogue between significantly different energy scales. Is it a coincidence that this phenomenon also appears in several mechanisms in the universe, which can be considered fundamental for good reasons? In three quite different mechanisms presented below, 1a) to 1c), such a delicate dependence of large length scales or small energy scales on the other end of the spectrum occurs. They have in common that physical explanations aiming to restore technical naturalness do not find application. On the other hand, the three successes of technical naturalness, 2a) to 2c), can be considered non-fundamental (as they are dealing with preliminary theories or composite entities). This comparison in 1) and 2) will provide the basis for hypothesis $H_{1}$.
  1. 1a)
    One sensitivity arises already in the mesocosm itself, without the need to know anything about quantum effects: in the theory of complex systems. This represents a significant phenomenological extension of linear Newtonian physics without leaving classical physics. In linear mechanics, the dynamics of a system is unaffected by infinitesimal disturbances. The dynamics of nonlinear systems, on the other hand, show a sensitive dependence on the initial conditions with the slightest influence - small changes have a significant effect (UV-IR mixing). This is particularly evident in the butterfly effect, provocatively expressed in the question of whether the flapping of a butterfly’s wings in Brazil can trigger a tornado in Texas (Lorenz 1972). More precisely, the presence of turbulence in a multitude of complex systems, extremely dependent on tiny perturbations of initial conditions, hints that a manifestation of broken autonomy of scales occurs as a fundamental property in the mesoscopic world as well, while the linear, i.e. effective theory describes only a small fraction of deterministic phenomena that show no dependencies on distant energy scales. Still, there are no discussions on physical explanations for broken autonomy of scales in nonlinear dynamics.
  2. 1b)